### v1

# Convert Language Model into a Value-based Strategic Planner

Abstract
Emotional support conversation (ESC) aims to alleviate the emotional distress of individuals through effective conversations. Although large language models (LLMs) have obtained remarkable progress on ESC, most of these studies might not define the diagram from the state model perspective, therefore providing a suboptimal solution for long-term satisfaction. To address such an issue, we leverage the Q-learning on LLMs, and propose a framework called straQ*. Our framework allows a plug-and-play LLM to bootstrap the planning during ESC, determine the optimal strategy based on long-term returns, and finally guide the LLM to response. Substantial experiments on ESC datasets suggest that straQ* outperforms many baselines, including direct inference, self-refine, chain of thought, finetuning, and finite state machines.
Convert Language Model into a Value-based Strategic Planner
Xiaoyu Wang1,2thanks: Work was done during the internship at Geely., Yue Zhao1, Qingqing Gu1, Zhonglin Jiang 1, Xiaokai Chen2, Yong Chen1, Luo Ji1thanks: Corresponding Author. 1 Geely AI Lab, Beijing, China 2 Beijing Institute of Technology, Beijing, China Correspondence: Luo.Ji1@geely.com
1 Introduction
Emotional Support Conversation (ESC) refers to dialogues aimed at alleviating a seeker’s emotional distress and challenges. Effective ESC is based on relational, psychological, and physical theories (Rains et al., 2020) and has been widely explored in artificial intelligence research Liu et al. (2021); Zhao et al. (2023). With advancements in LLMs, these models have shown strong performance in ESC Zheng et al. (2023); Kang et al. (2024). However, most LLM-based studies focus on immediate solutions without long-term support strategies. For example, while Liu et al. (2021) defines ESC in three stages (Exploration $\rightarrow$ Comforting $\rightarrow$ Action), LLMs often struggle with smooth transitions, leading to strategy biases.
Motivated by the recent progress of reinforcement learning (RL) on LLM-based studies Li et al. (2024b); Zhou et al. (2024); Wang et al. (2024a), we propose that ESC tasks can be defined as a strategy-level MDP, therefore value-based RL can help mitigate the aforementioned challenges. Given the current seeker’s utterance, emotion and conversational history, the LLM can be prompted to identify the long-term return of strategy, learn and produce the action value, and plan the optimal strategy. The determined strategy can then be prompted to another LLM to produce improved response, guided by the strategy.
Figure 1: Paradigm of straQ*.
In this paper, we propse a new framework called strategic Q* (straQ*), which converts LLM into a value-based strategic planner. We use the deep Q-learning (DQN) on LLM to provide a strategic Q function, with the strategy as the textual action. We use the averaged logits of actions to denote the Q value, and update the LLM parameter by the famous Bellman equation. By this manner, we convert the next-token prediction to next-strategy prediction, bootstrapping the TD loss of strategies instead of the original cross-entropy loss. This Q-net is used as a plug-and-play strategic planner, along with the conversation LLM produce the ultimate response. Our main contributions can be summarized as follows:
(1) We define the strategy-level MDP, and formulate the LLM architecture as a Q-function with textual input of state and strategy.
(2) We empirically verify that pretrained LLM can be finetuned by Bellman Equation and converges to optimum returns, with the averaged logit of action tokens as the q-value.
(3) Substantial experiments on ESConv and EmpatheticDialogues indicate that straQ* results in higher response quality and more reasonable planning of strategies.
(4) We design two reward machanisms including imitation and distillation, with the former better at automatic metrics, while the latter better at human scoring and generalization.
2 Preliminary
Strategy-level MDP.
In this work, we formalize the ESC tasks as an Markov decision process (MDP). We define the MDP $\mathcal{M}$ at the strategy level with $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$. We define the action space $a_{t}\in\mathcal{A}$ is the set of possible strategies, and $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the stepwise transition function.
Q-Learning.
In value-based RL, the goal is to learn both the state-value function $V(s)$ and the state-action value function $Q(s,a)$, therefore determining the optimal strategy from $Q$ also achieves the highest expected discounted cumulative reward:
$a^{\star}=$ $\arg\max_{a}Q(s,a)\leftarrow\arg\max\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$
which can be solved by the famous Bellman Equation:
$Q^{*}(s,a)=r(s,a)+\gamma\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime}).$ (1)
in which the superscript indicates the next step.
Instead of explicitly implementing the above equation, Deep Q-learning (DQN) approximates the maximization of the right-hand side with the deep value networks:
$\mathcal{L}(\theta)=|r(s,a)+Q_{\phi}(s^{\prime},a^{\prime})-Q_{\theta}(s,a)|^{2}$ (2)
where $\mathcal{L}$ is the loss, $\theta$ and $\phi$ are parameters of the Q-net and the targe Q-net, respectively. $\phi$ can be periodically synchronized from $\theta$.
Figure 2: Framework of straQ*.
3 Methodology
3.1 Task Definition
The problem of emotional-support conversation (ESC) can be characterized by an interleaved sequence of seeker $query$ and supporter $response$. To strengthen emotional-support performance, recent LM-based studies Liu et al. (2021) enhance the data content by augmenting the set of support strategies $\mathcal{A}$ and seeker emotions $\mathcal{E}$. For each conversation session, the background description can also be annotated on the session-level. Such augmented ESC can then be described as
$desc,\{query(t),e(t),a(t),resp(t)\}_{0:T}$ (3)
$a\in\mathcal{A},\quad e\in\mathcal{E}$
in which $T$ is the total number of conversation turn. At turn $t$, we denote the conversation history as
$h(t)=\{query,e(t),a(t),resp(t)\}_{0:T-1}$ (4)
Then the ESC sample at time $t$ can be alternatively expressed as $h(t),query(t),e(t),a(t),resp(t)$.
3.2 Framework
Our straQ* can be formally described as a 4-tuple $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T})$, where:
  • State. The state is a combination of description, emotion, history and query, i.e., $s=\{desc,e,h,query\}\in\mathcal{S}$.
  • Action. The strategy set is our action list.
  • Reward. An off-the-shelf foundation LLM generates the response upon on $a_{t}$, and the reward $r_{t}$ can be viewed as the instantaneous satisfaction of the seeker.
  • $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function. In the $t$-th turn, the seeker reacts further with a new query $q_{t+1}$ and emotion $e_{t+1}$. The history is updated by appending the current query and response, and the step is incremented by one.
3.3 Implementation on Language Models
LLM-based value function.
Our implementation starts from a pretrained LLM, with the parameter of $\theta$. We assume there is an instruction template with the placehoder of $s$, denoted by $\mathcal{I}(s)$. This instruction can be concatenated with $a$, $\mathcal{I}\oplus a$. Both state and state-action values can be obtained from the semantic understanding of LLM:
$Q_{\theta}(s,a)$ $\leftarrow\text{LLM}_{\phi}(\mathcal{I}(s)\oplus a)$ (5)
where $\leftarrow$ means to average the action logits.
Training a strategic value-function.
By replacing the Q-net in Equation 2 by the above expressions, we finetune the LLM by the Bellman Equation loss on last token logit. We keep the setting of the target Q-net, which is the same LLM architecture, while its parameter $\phi$ are periodically synchronized from $\theta$.
As in standard language modeling, the causal masking of the transformer allows us to perform Bellman updates on entire sequences in parallel.
Inference the optimal strategy.
Instead of decoding the next token, our finetuned LLM can produce the logits of available strategies, and the optimal strategy can be determined from the maximum logit
$a^{\star}\leftarrow\arg\max\text{LLM}(\mathcal{I}(s)\oplus a),a\in\mathcal{A}$ (6)
3.4 Definition of Rewards
Definition of rewards may be crucial when the action sampling is constrained by an offline dataset. In this paper, we study two reward mechanisms:
(1) Distillation: for each ($s$, $a$) pair from the dataset, we let a strong-basis LLM (e.g., GPT-4) to provide a judge score from 0 to 5. The detailed score prompt is in the Appendix. Since this manner distills the knowledge from a teacher model, we call this variant as straQ*-distill.
(2) Imitation: we consider each ($s$, $a$) pair from the dataset is an expert demonstration, therefore, always assigning a $+1$ reward. To amplify the distribution, we randomly sample different $a$ and assigning a $-1$ reward. The postive-negative ratio is 1:1. Since this manner belongs to the imitation learning, this variant is called straQ*-imit.
Figure 3: Training loss curves of straQ*
Methods Acc $\uparrow$ $\mathcal{Q}$uparrow $\mathcal{B}$downarrow B-2 $\uparrow$ R-L $\uparrow$
LLaMA3-8B-Instruct
+ Direct 11.80 10.26 1.61 3.47 10.64
+ Direct-Refine 17.08 11.07 1.27 3.10 6.13
+ Self-Refine 17.58 13.61 1.92 3.34 9.71
+ CoT 15.32 10.38 1.69 3.16 10.50
+ FSM 17.37 11.15 0.81 4.12 11.83
+ 1B straQ*-distill (ours) 41.22 38.95 0.57 3.89 11.80
+ 1B straQ*-imit (ours) 46.83 43.15 0.80 3.89 12.84
LLaMA3-8B-Instruct + SFT
+ Direct 32.43 21.29 1.28 6.97 16.59
+ CoT 30.80 17.70 1.35 6.51 15.00
+ FSM 28.83 18.36 1.32 7.57 17.42
+ 1B straQ*-distill (ours) 41.22 38.95 0.57 7.01 16.93
+ 1B straQ*-imit (ours) 46.83 43.15 0.80 7.63 17.30
Table 1: ID Results of automatic metrics including Acc, $\mathcal{Q}$, $\mathcal{B}$, Bleu-2 (B-2) and Rouge-L (R-L) on the testset of ESconv. The best results of each LLMs are bolded and the second best are underlined.
Methods B-2 R-L Dist-2 CIDEr
Direct 3.09 9.91 25.23 1.60
CoT 2.91 9.79 32.65 1.37
FSM 3.33 10.80 33.37 2.96
straQ*-distill (ours) 4.49 12.93 46.53 8.36
straQ*-imit (ours) 4.27 12.66 46.80 8.11
Table 2: OOD finetuned results of Bleu-2 (B-2) and Rouge-L (R-L) on EmpatheticDialogues. The best results of each LLMs are bolded and the second best are underlined.
Method Human Annotation
Fluency Emotion Acceptance Effectiveness Sensitivity Alignment Satisfaction
Orignal dataset 3.51 3.61 3.40 3.10 3.50 3.20 3.30
Llama3-8B-Instruct 2.95 3.00 2.60 2.40 2.70 2.70 2.60
+ Direct-Refine 3.09 3.09 2.73 2.91 2.91 2.82 2.84
+ Self-Refine 3.10 3.15 2.80 2.70 2.90 2.80 2.80
+ CoT 3.08 3.08 2.83 2.67 3.00 2.83 2.83
+ FSM 3.30 3.35 2.90 2.90 3.00 2.90 2.93
Llama3-8B-Instruct+ SFT 3.15 3.40 2.70 2.70 2.90 3.30 2.90
+ CoT 3.67 3.61 3.22 3.67 3.56 3.35 3.45
 + straQ*-distill (ours) 3.52 3.65 3.59 3.73 3.71 3.62 3.66
 + straQ*-imit (ours) 3.42 3.25 3.23 3.07 3.10 3.21 3.13
Table 3: Averaged Human evaluation of response quality on ESconv and EmpatheticDialogues.
4 Experiment
4.1 Setting
Implementation.
Llama3.2-1B-instruct AI@Meta (2024) is employed as the base model. Training is conducted on OpenRLHF Hu et al. (2024), with the learning rate of $65.0e-$ 6, window length of 2048 and epoch of 4. Batch size is set to 64. Experiments are conducted on 8 NVIDIA A100 80GB PCIe GPUs. In addition, the target network update frequency is set to 10, and the replay buffer size is 12,000.
Experiments.
Since our methodology is trained by annotated strategies, we employ ESconv as in-domain (ID) evaluation, and EmpatheticDialogues as the out-of-domain (OOD) evaluation, since EmpatheticDialogues does not have the strategy annotation. For the ID test, both strategy-related and response-related results can be provided. For the OOD test, we evaluate the zero-shot performance of the model finetuned on ESconv, with only response-related results provided.
Baselines.
We consider the following baselines:
(1) Direct: directly inference the LLM, with the same context.
(2) Direct-Refine: a straightforward refinement method in which the model revises its initial response to incorporate emotional support considerations.
(3) Self-Refine: a method (Madaan et al., 2023) initiates by generating feedback emphasizing emotional support from the initial response, then refining the response based on this feedback.
(4) CoT: uses the Chain-To-Thought prompt (Wei et al., 2022), which first generate the seeker’s emotion, which then guides the generation of strategy and response.
(5) FSM: the finite state machine Wang et al. (2024b) with finite sets of states and state-transitions triggered by inputs, and associated discrete actions.
4.2 Evaluation Methods
Automatic Metrics.
To evaluate the quality of strategy determination, we refer the evaluation methods proposed by Kang et al. (2024), which uses proficiency $\mathcal{Q}$ based on macro-F1 and the accuracy index (Acc) to represent the strategy accuracy, and preference bias $\mathcal{B}$ based on Bradley-Terry model (Bradley and Terry, 1952). Smaller $\mathcal{B}$ means less bias, therefore is better. For response quality, we utilize the famous Bleu-2 (B-2) and Rouge-L (R-L).
Human Scoring.
Similar with Kang et al. (2024), we annotate with the dimensions of Acceptance, Effectiveness, Sensitivity, Fluency, and Emotion, and the ultimate purpose, seeker’s Satisfaction. Scores are averaged over results of ESconv and EmpatheticDialogues.
4.3 Results
Training Curves.
Figure 3 shows the training loss curve of straQ*for 500 steps (approximately 3 epochs). Although the loss fluctuates significantly with the steps, it converges obviously overall, and finally tends to be stable.
Automatic Metrics.
Table 1 presents the evaluation of automatic indicators in domain (ID) across different base reply models. Compared to CoT and FSM, straQ*achieves higher strategy accuracy, lower bias, and greater similarity to source dataset replies. While straQ*-imit shows higher Acc due to dataset-derived rewards, this does not imply superiority over straQ*-distill, as discussed later. Since the EmpatheticDialogues dataset lacks strategy labels, OOD results in Table 2 focus on dialogue content, where straQ*demonstrates advantages and strong generalization. Details of per-strategy results are shown in the appendix.
Human Evaluation.
The results of the crowdsourcing evaluation shown in Table 3 indicate that straQ*-distill outperforms the baseline methods in various metrics, such as Fluency, Emotion, and Satisfaction. It also performs better than the replies in the source data. Conversely, straQ*-imit is slightly lower than the source data in performance. This is because the reward signal of straQ*-distill is the GPT-4 score, which prevents the generation of obviously poor responses present in the dataset.
Figure 4: Distribution of strategy determined by straQ*.
Left: the confusion matrix. Right: the transition matrix.
4.4 Discussion
Returns of Strategies.
Table 4 presents a statistical analysis of key performance indicators in reinforcement learning training. The <reward> represents GPT-4’s evaluation of generated replies on the test set, while the <value> reflects the model’s confidence in its decisions. Models trained with straQ*exhibit high <value>s, and straQ*-distill, leveraging GPT-4 scores as rewards, achieves a higher <reward>.
Method <reward> <value>
Original dataset 3.01 252.09
Llama3-8B-Instruct 3.66 346.31
straQ*-distill 3.99 424.78
straQ*-imit 3.72 445.95
Table 4: Average reward (eval by GPT-4) of strategy determination on the testset of ESConv.
Method Acc $\uparrow$ $\mathcal{Q}$uparrow $\mathcal{B}$downarrow B-2 $\uparrow$ R-L $\uparrow$
w/ value head 19.81 11.40 1.66 6.74 15.99
auto-regressive 46.22 43.01 0.69 7.25 16.48
straQ*-imit 46.83 43.15 0.80 7.63 17.30
Table 5: Ablation study of straQ*-imit on ESconv.
Figure 5: Sensitivity plots of straQ* on different $\gamma$ and model sizes. Metrics include Acc, $\mathcal{Q}$, B2 and R-L.
Strategy Prediction and Transitions.
The confusion matrix shown in Figure 4 indicates that for most strategies, the color is the darkest along the main diagonal. In the transition matrix, the color in the upper right triangular area is significantly darker than that in the lower left area, which is consistent with the three-stage theory in ESC.
Characters Strategy Response Reward Value
supporter - Hi! Hope you are doing well? 4 -
seeker - Hi, yes am doing well - -
supporter - How can I help you today? 3 -
seeker - I had a drinking challenge which has been affecting my marriage for quit sometime now - -
supporter Res.& Par. So as far as I understand that you had issues with your wife due to your drinking. Am I right? 3 7.6
straQ*-distill Res.& Par. It sounds like your drinking habits have been causing difficulties in your marriage for a while. 3 10.1
Direct Res.& Par. So, it sounds like there’s been a bit of tension in your marriage lately due to your drinking habits, and it even led to a confrontation that’s caused a rift between you and your wife. Can you tell me more about what’s been going on with your drinking? 4 8.0
seeker - Yes that is true - -
supporter Aff.& Rea. That is really a serious problem. I know it must be very tough time for you. Hope you are trying your best to save your marriage. 2 5.4
straQ*-distill Ref. That must be really difficult for you. I can imagine how much stress and worry this situation has brought into your life. 4 8.3
Direct Res.& Par. Just to make sure I understand correctly, it sounds like there was a confrontation with your wife about your drinking habits, and now you two aren’t speaking. Is that a fair summary of what happened? 3 4.7
seeker - Yes I have really reduced my drinking to ensure I save my marriage - -
supporter Ref. That’s is a great initiative. I know it is tough for you but not impossible. I can feel how hard you are trying t save your relationship. I appreciate you for that. 4 4.0
straQ*-distill Aff.& Rea. That’s a great step forward! It takes a lot of strength and commitment to make such a change, and you should be proud of yourself. 5 5.0
Direct Res.& Par. Just to make sure I understand correctly, you’ve been trying to cut down on your drinking to address your wife’s concerns, but despite your efforts, the issue still seems to be causing tension in your marriage, and now you’re not talking to each other. Is that a correct summary? 2 2.0
Table 6: Case of straQ* in ESconv.
4.5 Ablation Study
We conduct the following ablation studies:
  • w/ value head: Model is connected to a fully-connected network, and the output layer has only one neuron serving as the strategy value.
  • auto-regressive: Cross-entropy loss is used to calculate the loss instead of the reward signal.
Table 5 shows the straQ*model demonstrates better performance in various automatic metrics, which reflects the effectiveness of the proposed method.
4.6 Sensitivity Analysis
Figure 5 shows when $\gamma$ is within a certain range (0.8 to 0.9), its impact on the final performance is relatively small. However, when $\gamma$ is 0.99, the performance of the model decreases significantly. In addition, as the model size increases from 1B to 8B, the model performance also rises slowly. Nevertheless, considering the comprehensive factors of inference speed and computing resources, a 1B model is already enough.
4.7 Case
Table 6 presents the response scenarios of different models to the same set of multi-turn dialogues. straQ*-distill effectively achieves the three-stage strategy transfer and generally attains higher scores. In contrast, the strategy of the direct group remains unchanged throughout the multi-turn dialogues. As the number of turns increases, it becomes increasingly inappropriate, leading to lower scores.
5 Related Work
Typically, strategy learning occurs in goal-oriented conversations. Representative examples include DAT Li et al. (2024b). It represents conversations in a structured way by defining dialogue action tags and a multi-turn planner, and then generates responses based on these tags and pre-set planning strategies. In addition, DQ-HGAN Li et al. (2024a) realizes emotional support through the Graph Attention Network. ArCHer Zhou et al. (2024) proposes a hierarchical multi-turn RL algorithm to improve the efficiency and performance of language models in multi-turn goal-oriented decision-making tasks, and Q-star Wang et al. (2024a) improves multi-step reasoning for LLMs. However, such works usually do not have a clear explicit strategy, and their evaluation often relies on ground truth values.
In contrast, our straQ* learns to plan multi-turn explicit strategies offline from scoring feedback, and is decoupled from the dialogue model, which facilitates training and deployment..
6 Conclusion
In this paper, based on Q-learning, we propose a method named straQ*that can provide long-context emotional support conversation in multi-turn dialogue scenarios. We train to obtain the straQ*-imit model, which has a higher strategy accuracy and a higher response consistency on the Esconv dataset. In addition, we proposes the straQ*-distill method, which is trained from the reward signals of GPT-4’s scoring and achieves more excellent human evaluation results, and can be generalized to a broader range of scenarios.
7 Limitation
There are still some limitations of straQ*. The results of human evaluation may be biased, or deviate from the judgments of actual help - seekers due to the awareness of being engaged in scoring. Then, the testset may be small. Although it has little impact on the comparison between automated and human evaluations, sample sizes for some sub-categories may be insufficient when conducting a detailed analysis.
Appendix A Further Implementation Details
A.1 Dataset details
Emotional Support Conversations.
Emotional Support Conversation (ESC) is a task aimed at alleviating users’ negative emotions (e.g., anxiety, depression), where supporters assist seekers in managing emotions triggered by issues like work crises or interpersonal conflicts. Unlike emotion recognition tasks, ESC integrates psychological counseling mechanisms into the dialogue generation process, offering a deeper, context-sensitive solution for emotion regulation. In the practice, ESconv have exactly the aforementioned architecture while EmpatheticDialogues lacks the Strategy.
  • Emotion: Including emotion types and intensities, which help accurately capture the psychological state of the help-seeker.
  • Help-seeker profile: A brief survey before each conversation that provides insight into the current situation of the help-seeker, revealing the challenges they are facing.
  • Situation: A brief survey before each conversation that provides insight into the current situation of the help-seeker, revealing the challenges they are facing.
  • Strategy: The response rule selected for the current turn based on the seeker’s emotional state. There are eight predefined rules in total.
  • Response: The supporter’s response generated based on the history, inferred state, and selected rule.
ESconv.
Motivated by the Helping Skills Theory Hill (2009), Liu et al. (2021) divides ESC into three sequential stages: Exploration, Comforting, and Action, and proposes a dataset called ESconv. For each sample, the conversation is multi-turn, with the dialogue background and user emotion annotated. Upon each utterance of the supporter, 8 distinct support strategies are annotated. Table LABEL:tab:ex_ESConv exhibits an example of ESConv.
      Strategies Abbr. Definitions
Question Que. Inquiring about problem-related information to help the seeker clarify their issues, using open-ended questions for best results and closed questions for specific details.
Restatement or Paraphrasing Res.& Par. A simple, more concise rephrasing of the help-seeker’s statements that could help them see their situation more clearly.
Reflection of Feelings Ref. Articulate and describe the help-seeker’s feelings.
Self-disclosure Self-Dis. Divulge similar experiences that you have had or emotions that you share with the help-seeker to express your empathy.
Affirmation and Reassurance Aff.& Rea. Affirm the help seeker’s strengths, motivation, and capabilities and provide reassurance and encouragement.
Providing Suggestions Pro. Provide suggestions about how to change, but be careful to not overstep and tell them what to do.
Information Inf. Provide useful information to the help-seeker, for example with data, facts, opinions, resources, or by answering questions.
Others Others Exchange pleasantries and use other support strategies that do not fall into the above categories.
Table 7: Strategy definitions and Abbreviations of ESConv.
Emotion Type # Occurrence
anger 111
anxiety 354
depression 334
disgust 40
fear 95
nervousness 13
sadness 308
shame 42
Table 8: Emotion statistics of ESConv.
Table 7 provides definitions of support strategies in ESConv. Table 8 lists the emotion types and their occurrences in the dataset. The emotion types include anger, anxiety, depression, disgust, fear, nervousness, sadness, and shame. Table 10 summarizes the basic statistical information of ESConv.
EmpatheticDialogues.
EmpatheticDialogues is a dataset that consists of empathetic conversations. It aims to help in the development of empathetic language models by providing a large number of dialogues that express empathy.
A.2 Prompt format
To further strenthen the understanding capability of LLM on the strategy selection, we define the instruction as a multi-choice question (MCQ), forcing the LLM to choose one of the option numbers, instead of a plain question. Below is the content of the instruction template $\mathcal{I}$:
You are a psychological consultant providing support to a seeker. The seeker’s basic situation is as follows:
Emotion: {$e$}
Description: {$desp$}
Below is the conversation history between the seeker and the supporter:
{$h$}
The seeker’s current query is:
{$query$}
Based on the above context, please select the most appropriate response strategy from the following options:
strategy #(1) {$a_{1}$}
strategy #(k) {$a_{k}$}
Please provide your selection in the format of (1) through (k). Your selection is:
The prompt used in response generation:
You are a psychological consultant providing support to a seeker. The seeker’s basic situation is as follows:
Emotion: {$e$}
Description: {$desp$}
Below is the conversation history between the seeker and the supporter:
{$h$}
The seeker’s current query is:
{$query$}
The current response strategy is:
{$a$}
Based on the current response strategy and other information, please act as a supporter and provide the best response. Keep replies brief without additional pronouns or extra elements.
template of GPT-4 scoring.
You are a psychological consultant providing support to a seeker. The seeker’s basic situation is as follows:
Emotion: {$e$}
Description: {$desp$}
Below is the conversation history between the seeker and the supporter:
{$h$}
The seeker’s current query is:
{$query$}
Please evaluate whether the response is appropriate:
{$resp$}
Based on the information above, evaluate whether the response is suitable. Please remember to respond with a single integer number from 1 to 5, where 1 indicates "not suitable" and 5 indicates "very suitable". Please also provide a brief explanation of your decision.
Detailed results of Strategies.
The performance of the model often differs among various strategies. For humans, straightforward strategies like "Question" and "Providing Suggestions" may be more convenient to learn and apply. As a result, these strategies also appear relatively more frequently in the source data. Table 9 shows that for both the SFT and straQ*models, the above-mentioned strategies also have higher values of Acc, $\mathcal{Q}$, and R-L value. Moreover, compared with the SFT model, in addition to having higher indicators such as the average Acc, the straQ*model is more balanced among different strategies and avoids the occurrence of zero values.
Table 9: Per-strategy automatic metrics on ESconv.
Strategy Acc $\uparrow$ $\mathcal{Q}$uparrow $\mathcal{B}$downarrow B-2 $\uparrow$ R-L $\uparrow$ Dist-2 CIDEr
SFT Que. 57.52 48.24 1.60 9.37 21.88 64.42 34.22
Res.& Par. 18.52 20.55 0.84 7.96 16.99 77.38 21.12
Ref. 1.57 2.92 0.08 5.74 14.90 73.44 11.72
Self-Dis. 2.36 4.38 0.06 4.99 12.30 76.71 7.85
Aff.& Rea. 20.09 22.45 0.83 5.94 15.17 70.72 13.20
Pro. 75.22 40.95 4.12 5.99 14.26 71.26 11.04
Inf. 0.00 0.00 0.00 5.93 12.24 78.54 12.95
Others 23.20 30.77 0.46 8.38 18.21 74.08 27.78
straQ*-distill Que. 71.07 60.59 2.43 9.48 22.05 64.44 33.50
Res.& Par. 8.97 16.67 0.16 8.45 17.30 79.05 21.60
Ref. 40.88 38.69 0.43 5.18 13.62 75.95 9.78
Self-Dis. 29.85 41.75 0.46 4.90 12.70 76.00 6.01
Aff.& Rea. 36.63 42.11 0.47 6.30 15.46 70.35 13.66
Pro. 69.08 56.13 0.60 5.95 14.03 70.51 10.62
Inf. 26.67 40.14 0.47 5.33 13.21 79.94 8.11
Others 45.93 45.95 0.53 6.60 15.19 71.85 24.34
Category ESconv EmpatheticDialogues
# Sessions 1.3K 2.5K
# Uttr 38K 11.0K
Average # Uttr 28.9 4.3
Average Uttr Len 18.8 16.7
Seeker # Uttr 20K 5.7K
Avg # Uttr 15.4 2.2
Avg Uttr Len 16.8 20.8
# Emotions 11 32
Supporter # Uttr 18K 5.2K
Avg # Uttr 13.6 2.1
Avg Uttr Len 21.0 12.3
# Strategies 8 -
Table 10: Statistics of ESConv and EmpatheticDialogues. ‘Uttr’ abbreviates Utterance.
We start with the criteria proposed by Kang et al. (2024). The human evaluation is aimed to algin with the ultimate purpose of ESC, the seeker’s satisfaction. To achieve this, the supporter’s behavior can be further classified into the following criteria:
Acceptance: Does the seeker accept without discomfort;
Effectiveness: Is it helpful in shifting negative emotions or attitudes towards a positive direction;
Sensitivity: Does it take into consideration the general state of the seeker. Furthermore, to clarify the capability of LLMs to align strategy and responses, we include Alignment.
To achieve a more elaborate assessment, we consider three more dimensions addressing the generation quality:
Fluency: the level of fluency of response.
Emotion: the emotional intensity of response which could affect the seeker’s emotion state.
Interesting: Whether the response can arouse the seeker’s interest and curiosity, presenting unique ideas, vivid expressions or engaging elements that capture the seeker’s attention and make the interaction more appealing.
We engage our interns as human evaluators to rate the models according to these multiple aspects, namely Fluency, Emotion, Interesting, and Satisfaction, with Satisfaction covering Acceptance, Effective, Sensitivity, and Satisfaction itself.
Throughout this evaluation process, we strictly comply with international regulations and ethical norms, ensuring that all practices conform to the necessary guidelines regarding participant involvement and data integrity.
To guarantee the accuracy and reliability of the evaluation results, a pre - evaluation training program is meticulously designed and implemented. During this training, the evaluation criteria are clearly and systematically expounded. Moreover, detailed explanations and scoring rules corresponding to each score are provided.
Evaluators are required to independently evaluate each sample in strict accordance with the pre - established criteria. By adhering to these principles, the evaluation process maintains objectivity, standardization, and consistency, thus enhancing the overall quality and credibility of the evaluation results.
The detailed manual scoring criteria are as follows:
  • Fluency:
    1: The sentence is highly incoherent, making it extremely difficult to understand and failing to convey a meaningful idea.
    2: The sentence has significant incoherence issues, with only parts of it making sense and struggling to form a complete thought.
    3: The sentence contains some incoherence and occasional errors, but can still convey the general meaning to a certain extent.
    4: The sentence is mostly fluent with only minor errors or slight awkwardness in expression, and effectively communicates the intended meaning.
    5: Perfect. The sentence is completely fluent, free of any errors in grammar, punctuation, or expression, and clearly conveys the idea.
  • Emotion:
    1: The emotional expression is extremely inappropriate and chaotic, not in line with the content, and may convey wrong emotions.
    2: The emotional expression has obvious flaws, either too weak or exaggerated, and is disjointed from the content.
    3: The emotional expression is average. It can convey basic emotions but lacks depth and has minor issues.
    4: The emotional expression is good. It can effectively convey the intended emotion with an appropriate intensity and is well integrated with the content.
    5: The emotional expression is excellent. It is rich, nuanced, and perfectly matches the content, capable of evoking a strong and appropriate emotional response.
  • Acceptance:
    1: The response inescapably triggers emotional resistance.
    2: The response is highly likely to trigger emotional resistance.
    3: The response has a possibility of emotional resistance occurring.
    4: The response rarely provokes emotional resistance.
    5: The response has no occurrence of emotional resistance.
  • Effectiveness:
    1: The response actually worsens the seeker’s emotional distress.
    2: The response carries the risk of increasing stress levels, and this outcome varies depending on the individual user.
    3: The response fails to alter the seeker’s current emotional intensity and keeps it at the same level.
    4: The response shows promise in calming the emotional intensity; however, it is overly complicated or ambiguous for the user to fully comprehend and utilize effectively.
    5: The response appears to be highly effective in soothing the seeker’s emotions and offers valuable and practical emotional support.
  • Sensitivity:
    1: The response renders inaccurate evaluations regarding the seeker’s state.
    2: The response is characterized by rash judgments, as it lacks adequate assessment and in-depth exploration of the seeker’s state.
    3: The response is formulated with a one-sided judgment and a limited exploration of the seeker’s state.
    4: The response demonstrates an understanding that only covers a part of the seeker’s state.
    5: The response precisely grasps the seeker’s state and is appropriately tailored according to the seeker’s actual situation.
  • Alignment:
    1: The response is in total contradiction to the predicted strategy.
    2: The response has a minor deviation from the predicted strategy.
    3: There is some ambiguity between the response and the predicted strategy.
    4: The response largely matches the predicted strategy, yet it contains some ambiguous elements.
    5: The response effectively makes itself consistent with the predicted strategy.
  • Satisfaction:
    1: The response is extremely disappointing. It doesn’t answer the question at all and is of no help.
    2: The response is poor. It only gives a partial answer and leaves many doubts unresolved.
    3: The response is average. It meets the basic requirements but isn’t particularly outstanding.
    4: The response is good. It answers the question clearly and provides some useful details.
    5: The response is excellent. It not only answers the question perfectly but also offers valuable additional insights.
Strategy Count Ratio Max Min Avg Median
Que. 2574 17.6% 5 1 3.54 4
Res.& Par. 981 6.7% 5 1 3.48 3
Ref. 1253 8.6% 5 2 3.65 4
Self-Dis. 1410 9.6% 5 2 3.48 3
Aff.& Rea. 2444 16.7% 5 1 3.76 4
Pro. 2367 16.2% 5 1 3.77 4
Inf. 995 6.8% 5 2 3.75 4
Others 2600 17.8% 5 1 3.77 4
Total 14624 100.0% 5 1 3.67 4
Table 11: Statistics of GPT-4 score.