### v3

# Convert Language Model into a Value-based Strategic Planner

Abstract
Emotional support conversation (ESC) aims to alleviate the emotional distress of individuals through effective conversations. Although large language models (LLMs) have obtained remarkable progress on ESC, most of these studies might not define the diagram from the state model perspective, therefore providing a suboptimal solution for long-term satisfaction. To address such an issue, we leverage the Q-learning on LLMs, and propose a framework called straQ*. Our framework allows a plug-and-play LLM to bootstrap the planning during ESC, determine the optimal strategy based on long-term returns, and finally guide the LLM to response. Substantial experiments on ESC datasets suggest that straQ* outperforms many baselines, including direct inference, self-refine, chain of thought, finetuning, and finite state machines.
Convert Language Model into a Value-based Strategic Planner
Xiaoyu Wang1,2thanks: Work was done during the internship at Geely., Yue Zhao1, Qingqing Gu1, Zhonglin Jiang 1, Xiaokai Chen2, Yong Chen1, Luo Ji1thanks: Corresponding Author. 1 Geely AI Lab, Beijing, China 2 Beijing Institute of Technology, Beijing, China Correspondence: Luo.Ji1@geely.com
1 Introduction
Emotional Support Conversation (ESC) refers to dialogues aimed at alleviating a seeker’s emotional distress and challenges. Effective ESC is based on relational, psychological, and physical theories (Rains et al., 2020) and has been widely explored in artificial intelligence research Liu et al. (2021); Zhao et al. (2023). With advancements in LLMs, these models have shown strong performance in ESC Zheng et al. (2023); Kang et al. (2024). However, most LLM-based studies focus on immediate solutions without long-term support strategies. For example, while Liu et al. (2021) defines ESC in three stages (Exploration $\rightarrow$ Comforting $\rightarrow$ Action), LLMs often struggle with smooth transitions, leading to strategy biases.
Motivated by the recent progress of reinforcement learning (RL) on LLM-based studies Li et al. (2024b); Zhou et al. (2024); Wang et al. (2024a), we propose that ESC tasks can be defined as a strategy-level MDP, therefore value-based RL can help mitigate the aforementioned challenges. Given the current seeker’s utterance, emotion and conversational history, the LLM can be prompted to identify the long-term return of strategy, learn and produce the action value, and plan the optimal strategy. The determined strategy can then be prompted to another LLM to produce an improved response, guided by the strategy.
Figure 1: Paradigm of straQ*. A plug-and-play LLM-based planner selects the optimal strategy from maximized $Q$, then steers the LLM to enhance the response.
In this paper, we propose a new framework called strategic Q* (straQ*), which converts LLM into a value-based strategic planner. We use deep Q-learning (DQN) on LLM to provide a strategic Q function, with the strategy as the textual action. We use the averaged logits of actions to denote the Q value, and update the LLM parameter by the famous Bellman equation. By this manner, we convert the next-token prediction to next-strategy prediction, bootstrapping the TD loss of strategies instead of the original cross-entropy loss. This Q-net is used as a plug-and-play strategic planner, along with the conversation LLM to produce the ultimate response. Our main contributions can be summarized as follows:
(1) We define the strategy-level MDP, and formulate the LLM architecture as a Q-function with textual input of state and strategy.
(2) We empirically verify that pretrained LLM can be finetuned by the Bellman Equation and converges to optimum returns, with the averaged logit of action tokens as the q-value.
(3) Substantial experiments on ESConv and EmpatheticDialogues indicate that straQ* results in higher response quality and more reasonable planning of strategies.
(4) We design two reward mechanisms including imitation and distillation, with the former better at automatic metrics, while the latter better at human scoring and generalization.
2 Preliminary
Strategy-level MDP.
The Markov decision process (MDP) is usually defined as a 5-tuple $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},\gamma)$, where $\mathcal{S}$ is the state set, $\mathcal{A}$ is the action set, $\mathcal{R}$ is the reward set, $\gamma$ is the discounting factor of rewards, and $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the state transition function. In this work, we formalize the ESC task as a strategy-level MDP, with the action space $\mathcal{A}=\{a\}$ as the set of possible strategies.
Q-Learning.
In value-based RL, the goal is to learn the state-value function $V(s)$ or the state-action value function $Q(s,a)$, such that the determined action achieves the highest expected discounted cumulative reward:
$a^{\star}=$ $\arg\max_{a}Q(s,a)\leftarrow\arg\max\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$
which is solved by the famous Bellman Equation:
$Q^{*}(s,a)=r(s,a)+\gamma\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})$ (1)
in which the superscript indicates the next step. Instead of explicitly implementing the above equation, Deep Q-learning (DQN) approximates the maximization of the right-hand side with the deep value networks:
$\mathcal{L}(\theta)=|r(s,a)+Q_{\phi}(s^{\prime},a^{\prime})-Q_{\theta}(s,a)|^{2}$ (2)
where $\mathcal{L}$ is the loss, $\theta$ and $\phi$ are parameters of the Q-net and the target Q-net, respectively. $\phi$ can be periodically synchronized from $\theta$.
Figure 2: The training framework of straQ*. Averaged log probability of action tokens is defined as $Q(s,a)$ which deduces the training loss from Bellman Equation.
3 Methodology
3.1 Task Definition
The problem of emotional-support conversation (ESC) can be characterized by an interleaved sequence of seeker $query$ and supporter $response$. To strengthen emotional-support performance, recent studies Liu et al. (2021); Rashkin et al. (2019) enhance the data content by augmenting the set of support strategies $\mathcal{A}$ and seeker emotions $\mathcal{E}$. For each conversation session, the background $description$ is also annotated on the session-level. Such augmented ESC can then be described as
$desc,\{query(t),e(t),a(t),resp(t)\}_{0:T}$ (3)
$a\in\mathcal{A},\quad e\in\mathcal{E}$
in which $desp$ and $resp$ are the abbreviations of $descrition$ and $response$, and $T$ is the total number of conversation turns. At turn $t$, we denote the conversation history as
$h(t)=\{query(t),e(t),a(t),resp(t)\}_{0:T-1}$ (4)
Then the ESC sample at time $t$ can be alternatively expressed as $\{h(t),query(t),e(t),a(t),resp(t)\}$.
3.2 System Variables
We define important system variables as follows:
$\bullet$ State: The state is a combination of description, emotion, history and query, i.e., $s=\{desc,e,h,query\}\in\mathcal{S}$.
$\bullet$ Action: The conversational strategy, $a\in\mathcal{A}$.
$\bullet$ Reward: The reward $r_{t}$ can be viewed as the instantaneous satisfaction of the seeker, which can be either inferred from an annotated datasets, or generated by an off-the-shelf model evaluator (LLM-as-the-Judge).
$\bullet$ $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function. After the $t$-th turn, $h$ is updated by appending the current $query(t)$ and $resp(t)$, the seeker reacts further with new $query(t+1)$ and $e(t+1)$, and the step is incremented by one.
3.3 Implementation on Language Models
LLM-based value function.
Our implementation starts from a pretrained LLM, with the parameter of $\theta$. We assume there is an instruction template with the placeholder of $s$, denoted by $\mathcal{I}(s)$. This instruction can be concatenated with $a$, $\mathcal{I}\oplus a$. Both state and state-action values can be obtained from the semantic understanding of LLM:
$Q_{\theta}(s,a)$ $\leftarrowtail\text{LLM}_{\theta}(\mathcal{I}(s)\oplus a)$ (5)
[…]
Training a strategic value-function.
By replacing the Q-net in Equation 2 by the above expressions, we finetune the LLM by the Bellman Equation loss on last token logit. As in standard language modeling, the causal masking of the transformer allows us to perform Bellman updates on entire sequences in parallel. Figure 2 exhibits this training framework.
We keep the setting of the target Q-net, which is the same LLM architecture, while its parameter $\phi$ is periodically synchronized from $\theta$.
Inference the optimal strategy.
Instead of decoding the next token, the finetuned LLM produces logits of available strategies, and the optimal strategy can be determined from the maximum logit
$a^{\star}\leftarrow\arg\max\text{LLM}(\mathcal{I}(s)\oplus a),a\in\mathcal{A}$ (6)
Figure 3: Diagram of our pipeline. Reward is annotated by the imitation or distillation strategy. The action is chosen from the strategy which maximizes the $Q$ value.
Instruction template.
We briefly exhibit our instruction $\mathcal{I}(s)$ here:
[…]
with the full version in Appendix B. By forming the prompt To further strengthen the understanding capability of LLM on the strategy selection, we formulate $\mathcal{I}$ as a multi-choice question (MCQ), instead of a plain question, forcing the LLM to choose one of the option numbers. Accordingly, the action set becomes the set of possible strategy index $a\in\mathcal{A}:=\{1,2,\cdots,K\}$ where $K$ is the total number of strategies.
3.4 Reward Definitions
Choice of rewards may be crucial especially when the sampling is constrained by an offline dataset. In this paper, we study two reward mechanisms:
(1) Distillation: for each ($s$, $a$) pair from the dataset, we let a strong-basis LLM (e.g., GPT-4) to provide a judge score from 0 to 5 (The detailed judge prompt is in Appendix B). Since this manner distills the knowledge from a teacher model, we call this variant as straQ*-distill.
(2) Imitation: we consider each ($s$, $a$) pair from the dataset is an expert demonstration, therefore, always assigned with $r$ of $+1$. To amplify the distribution, we randomly sample a different $a$ and assign with $r$ of $-1$. The positive-negative ratio is 1:1. Since this manner imitates the positive samples directly, this variant is called straQ*-imit.
Figure 3 shows the entire pipeline of straQ*.
Figure 4: Training loss curve of straQ*.
Strategies Abbr. Stage
Question Que. I
Restatement or Paraphrasing Res.& Par. I
Reflection of Feelings Ref. II
Self-disclosure Self-Dis. II
Affirmation and Reassurance Aff.& Rea. III
Providing Suggestions Pro. III
Information Inf. III
Others Others -
Table 1: Strategy names, abbreviations and stages.
4 Experiment
4.1 Setting
Implementation.
Llama3.2-1B-instruct AI@Meta (2024) is employed as the base model. Training is conducted on OpenRLHF Hu et al. (2024), with the learning rate of $65.0e-$ 6, window length of 2048, batch size of 64, and epoch of 4. The target network update frequency is set to 10, the replay buffer size is 12,000, and $\gamma=0.85$.
Datasets.
Training of straQ* requires the annotation of strategies. We use ESConv Liu et al. (2021) as the training set and also the in-domain (ID) test. ESConv provides $K=8$ strategies, which belong to three ESC stages: Exploration (I), Comforting (II) and Action (III). Table 1 shows their full names, abbreviations and corresponding stages.
Furthermore, EmpatheticDialogues Rashkin et al. (2019) is employed as the out-of-domain (OOD) evaluation, since EmpatheticDialogues does not have the strategy annotation. For the ID test, both strategy-related and response-related results can be provided. For the OOD test, only zero-shot response-related results are provided. Appendix A provides a more detailed introduction of ESConv and EmpatheticDialogues.
4.2 Evaluation Methods
Automatic Metrics.
To evaluate the quality of strategy determination, we refer the evaluation methods proposed by Kang et al. (2024), which uses proficiency $\mathcal{Q}$ based on macro-F1, and preference bias $\mathcal{B}$ based on Bradley-Terry model (Bradley and Terry, 1952). Smaller $\mathcal{B}$ means less bias, therefore is better. We also include the strategy prediction accuracy (Acc). For response quality, we utilize the famous Bleu-2 (B-2), Rouge-L (R-L), Distinct-2 (D-2) and CIDEr, calculating from the similarity with the ground truth response.
Human Scoring.
Similar with Kang et al. (2024), we annotate with the dimensions of Acceptance, Effectiveness, Sensitivity, Fluency, and Emotion, and the ultimate purpose, seeker’s Satisfaction.
Baselines.
We consider the following baselines:
(1) Direct: directly inference the LLM.
(2) Direct-Refine: the model immediately revises its response within the same utterance to incorporate emotional support considerations.
(3) Self-Refine (Madaan et al., 2023): the model considers the emotional support, generates a feedback from the initial response, then refines the response based on the feedback.
(4) CoT (Wei et al., 2022): steered by the chain-of-thought prompt, the model first identifies emotion, then generates strategy, and finally response.
(5) FSM Wang et al. (2024b): the finite state machine with finite sets of states and state-transitions triggered by inputs, and associated discrete actions.
Methods Acc $\uparrow$ $\mathcal{Q}$uparrow $\mathcal{B}$downarrow B-2 $\uparrow$ R-L $\uparrow$
LLaMA3-8B-Instruct
Direct 11.80 10.26 1.61 3.47 10.64
+ Direct-Refine 17.08 11.07 1.27 3.10 6.13
+ Self-Refine 17.58 13.61 1.92 3.34 9.71
+ CoT 15.32 10.38 1.69 3.16 10.50
+ FSM 17.37 11.15 0.81 4.12 11.83
+ 1B straQ*-distill (ours) 41.22 38.95 0.57 3.89 11.80
+ 1B straQ*-imit (ours) 46.83 43.15 0.80 3.89 12.84
LLaMA3-8B-Instruct + SFT
Direct 32.43 21.29 1.28 6.97 16.59
+ CoT 30.80 17.70 1.35 6.51 15.00
+ FSM 28.83 18.36 1.32 7.57 17.42
+ 1B straQ*-distill (ours) 41.22 38.95 0.57 7.01 16.93
+ 1B straQ*-imit (ours) 46.83 43.15 0.80 7.63 17.30
Table 2: ID Results of automatic metrics including Acc, $\mathcal{Q}$, $\mathcal{B}$, Bleu-2 (B-2) and Rouge-L (R-L) on the testset of ESConv. The best results of each LLMs are bolded and the second best are underlined.
Methods B-2 R-L Dist-2 CIDEr
Direct 3.09 9.91 25.23 1.60
+ CoT 2.91 9.79 32.65 1.37
+ FSM 3.33 10.80 33.37 2.96
+ 1B straQ*-distill (ours) 4.49 12.93 46.53 8.36
+ 1B straQ*-imit (ours) 4.27 12.66 46.80 8.11
Table 3: OOD finetuned results of Bleu-2 (B-2) and Rouge-L (R-L) on EmpatheticDialogues. The best results of each LLMs are bolded and the second best are underlined.
Method Human Annotation
Fluency Emotion Acceptance Effectiveness Sensitivity Alignment Satisfaction
Original dataset 3.51 3.61 3.40 3.10 3.50 3.20 3.30
Llama3-8B-Instruct 2.95 3.00 2.60 2.40 2.70 2.70 2.60
+ Direct-Refine 3.09 3.09 2.73 2.91 2.91 2.82 2.84
+ Self-Refine 3.10 3.15 2.80 2.70 2.90 2.80 2.80
+ CoT 3.08 3.08 2.83 2.67 3.00 2.83 2.83
+ FSM 3.30 3.35 2.90 2.90 3.00 2.90 2.93
Llama3-8B-Instruct+ SFT 3.15 3.40 2.70 2.70 2.90 3.30 2.90
+ CoT 3.67 3.61 3.22 3.67 3.56 3.35 3.45
 + straQ*-distill (ours) 3.52 3.65 3.59 3.73 3.71 3.62 3.66
 + straQ*-imit (ours) 3.42 3.25 3.23 3.07 3.10 3.21 3.13
Table 4: Averaged Human evaluation of response quality on ESConv and EmpatheticDialogues.
4.3 Results
Training Curves.
Figure 4 shows the training loss curve of straQ*for 500 steps (approximately 3 epochs). Although the loss initially fluctuates significantly, it adapts to the new training paradigm, and finally tends to be stable.
Automatic Evaluations.
Table 2 presents the automatic metrics on the ID evaluation, with the basis of either the original LLM, or the specifically finetuned version. Compared to baselines, straQ*generally achieves higher strategy accuracy, lower bias, and higher similarity to the ground truth responses. Furthermore, straQ*-imit performs better than straQ*-distill on this setting, suggesting that the imitation-version of rewards result in better ID performance.
In Table 3, we further compare the OOD results of the models finetuned by ESConv, with the strategy lists inferred from ESConv. Results suggest that straQ*demonstrates strong generalization than these baselines. Specifically, straQ*-distill surpasses straQ*-imit this time, indicating the distilled knowledge from the teacher model is more general than simply imitating a limited dataset.
Human Evaluation.
The results of the crowdsourcing evaluation shown in Table 4 indicate that straQ*-distill outperforms the baseline methods in various metrics, such as Fluency, Emotion, and Satisfaction. It also performs better than the replies in the source data. Conversely, straQ*-imit is slightly lower than the source data in performance. Using the GPT-4 score as a reward, straQ* can determine strategies more from the aspect of performance optimization, not simply imitating the demonstration.
Figure 5: Distribution of strategy determined by straQ*. Strategies are labeled with the stage index (I, II, III) which represents the general scenario Exploration $\rightarrow$ Comforting $\rightarrow$ Action in ESC. Left: the confusion matrix (acted strategy (row) VS ground truth strategy (column)). Right: the transition matrix (acted strategy (row) VS the next-acted strategy (column)).
Ablation Study.
Two ablations are studied:
(1) w/ value head: append the model with a classification head which produces the score logit.
(2) auto-regressive: keep the cross-entropy loss with the ground truth action as the target text.
In more detail, w/ value head is the usual solution for a reward model in RLHF, while auto-regressive can be viewed as a standard fine-tuning solution for the strategic planner. Table 6 shows that straQ*outperforms both of them in various automatic metrics, indicating our methodology can better align with the strategy semantics and more accurately capture the strategic value.
Sensitivity Analysis.
Figure 6 shows the ID performance evolutions on different $\gamma$ choices. Smaller $\gamma$ means we are more focused on the transient performance and relatively neglect the long-term value. Results show that the optimal accuracy of strategy happens on $\gamma=0.9$, while the best response-related metrics correspond to $\gamma=0.85$. Because B-2 and R-L are similarity-based, the current reward is more relative to them than future rewards. Therefore, this observation is reasonable.
4.4 Discussions
Scalability and application.
Figure 6 (bottom-right) also compares the B-2 results on different model sizes. As the model becomes larger, the performance also increases, indicating straQ* can have good scalability. However, larger models result in higher computation overhead and slower speed, which may hinder the practical application of straQ*. Therefore, in the formal application, we still adhere to the 1B choice, employing it as a lightweight planner.
From previous results, we utilize straQ*-distill in the actual application to have better generalization and better alignment with human knowledge.
Returns of Strategies.
Table 5 further analyzes two important indicators of value-based RL, the averaged rewards and values. In this analysis, the rewards are provided by GPT-4. straQ* achieve both higher <reward> and <value> than direct inference of the base model, as well as the annotation of the original dataset. This result shows that straQ* statistically obtains higher returns, which is the primary purpose of Q-Learning.
Method <reward> <value>
Original dataset 3.01 252.09
Llama3-8B-Instruct 3.66 346.31
straQ*-distill (ours) 3.99 424.78
straQ*-imit (ours) 3.72 445.95
Table 5: Average reward (eval by GPT-4) of strategy determination on the test set of ESConv.
Method Acc $\uparrow$ $\mathcal{Q}$uparrow $\mathcal{B}$downarrow B-2 $\uparrow$ R-L $\uparrow$
w/ value head 19.81 11.40 1.66 6.74 15.99
auto-regressive 46.22 43.01 0.69 7.25 16.48
straQ*-imit 46.83 43.15 0.80 7.63 17.03
Table 6: Ablation study of straQ*-imit on ESConv.
Figure 6: Sensitivity plots of straQ*-imit on different $\gamma$ and model sizes. Metrics include Acc, $\mathcal{Q}$, B2, and R-L.
Strategy Prediction and Transitions.
Figure 5 (Left) exhibits the confusion matrix of strategies, with the rows representing the prediction, and the columns representing the ground truth. Results show that most occurrences happen on the diagonal grids, verifying the prediction accuracy.
Figure 5 (Right) visualizes the transition matrix. A grid $(i,j)$ means the strategy $i$ to the strategy $j$, where the strategies are sorted from their ESC stages (from I to III) for both rows and columns. Therefore, transitions from an early stage to a later stage should occur on the upper-triangle region of the transition matrix. The results in the figure validate this proposition.
Characters Strategy Response Reward Value
supporter - Hi! Hope you are doing well? - -
seeker - Hi, yes am doing well - -
supporter - How can I help you today? - -
seeker - I had a drinking challenge which has been affecting my marriage for quit sometime now - -
supporter Res.& Par. (I) So as far as I understand that you had issues with your wife due to your drinking. Am I right? 3 7.6
Direct Res.& Par. (I) So, it sounds like there’s been a bit of tension in your marriage lately due to your drinking habits, and it even led to a confrontation that’s caused a rift between you and your wife. Can you tell me more about what’s been going on with your drinking? 4 8.0
straQ*-distill Res.& Par. (I) It sounds like your drinking habits have been causing difficulties in your marriage for a while. 3 10.1
seeker - Yes that is true - -
supporter Aff.& Rea. (II) That is really a serious problem. I know it must be very tough time for you. Hope you are trying your best to save your marriage. 2 5.4
Direct Res.& Par. (I) Just to make sure I understand correctly, it sounds like there was a confrontation with your wife about your drinking habits, and now you two aren’t speaking. Is that a fair summary of what happened? 3 4.7
straQ*-distill Ref. (II) That must be really difficult for you. I can imagine how much stress and worry this situation has brought into your life. 4 8.3
seeker - Yes I have really reduced my drinking to ensure I save my marriage - -
supporter Ref. (II) That’s is a great initiative. I know it is tough for you but not impossible. I can feel how hard you are trying t save your relationship. I appreciate you for that. 4 4.0
Direct Res.& Par. (I) Just to make sure I understand correctly, you’ve been trying to cut down on your drinking to address your wife’s concerns, but despite your efforts, the issue still seems to be causing tension in your marriage, and now you’re not talking to each other. Is that a correct summary? 2 2.0
straQ*-distill Aff.& Rea. (III) That’s a great step forward! It takes a lot of strength and commitment to make such a change, and you should be proud of yourself. 5 5.0
Table 7: Case of straQ* in ESConv.
Detailed results of Strategies.
Strategies’ popularities and occurrences may differ in nature. For example, straightforward strategies like "Question" and "Providing Suggestions" may be more convenient to learn and apply. A reasonable strategic planner should reflect this frequency difference, but also avoid significant bias (e.g., always determines the most frequent strategy). To further study the strategy-related performance, we further exhibit the per-strategy results in Table 8, with the comparison between straQ*-distill and SFT. One can observe that straQ* model is more balanced across different strategies, and SFT even has zero predictions on the strategy of "Information".
Table 8: Per-strategy automatic metrics on ESConv.
Strategy Acc $\uparrow$ $\mathcal{Q}$uparrow $\mathcal{B}$downarrow B-2 $\uparrow$ R-L $\uparrow$ Dist-2 CIDEr
SFT Que. 57.52 48.24 1.60 9.37 21.88 64.42 34.22
Res.& Par. 18.52 20.55 0.84 7.96 16.99 77.38 21.12
Ref. 1.57 2.92 0.08 5.74 14.90 73.44 11.72
Self-Dis. 2.36 4.38 0.06 4.99 12.30 76.71 7.85
Aff.& Rea. 20.09 22.45 0.83 5.94 15.17 70.72 13.20
Pro. 75.22 40.95 4.12 5.99 14.26 71.26 11.04
Inf. 0.00 0.00 0.00 5.93 12.24 78.54 12.95
Others 23.20 30.77 0.46 8.38 18.21 74.08 27.78
straQ*-distill Que. 71.07 60.59 2.43 9.48 22.05 64.44 33.50
Res.& Par. 8.97 16.67 0.16 8.45 17.30 79.05 21.60
Ref. 40.88 38.69 0.43 5.18 13.62 75.95 9.78
Self-Dis. 29.85 41.75 0.46 4.90 12.70 76.00 6.01
Aff.& Rea. 36.63 42.11 0.47 6.30 15.46 70.35 13.66
Pro. 69.08 56.13 0.60 5.95 14.03 70.51 10.62
Inf. 26.67 40.14 0.47 5.33 13.21 79.94 8.11
Others 45.93 45.95 0.53 6.60 15.19 71.85 24.34
Typical Case.
Table 7 presents a typical case of straQ*-distill in the third turn, comparing to Direct, and also the original response in the dataset. To better illustrate the effectiveness of the value learning, we present the reward and value scores for each response. In this case, straQ* does not simply maximize the immediate reward, but maximizes the long-term return (i.e., the value), which is calculated from subsequent turns. Also, straQ* in this case exhibits a perfect stage-turnover, guiding the conversation from the first stage (strategy Res.&Par.), then the second stage (strategy Ref.), to the third stage (strategy Aff.& Rea.). Comparing to Direct (stays in I) the original response (I to II to II), planning of straQ* is more consistent with the theory proposed in Liu et al. (2021).
5 Related Work
There are some RL studies in goal-oriented conversations Li et al. (2024b); Zhou et al. (2024); Li et al. (2024a). For example, DAT Li et al. (2024b) defines dialogue action tags, and then generates responses by multi-turn planning. ArCHer Zhou et al. (2024) proposes a hierarchical RL algorithm to improve the efficiency and performance of LLMs. These works adapt the conversational LLM, and rely on ground truth annotations. In contrast, our straQ* implements an explicit, lightweight and plug-and-play planner, which balances the foundation capability and the strategic thinking.
6 Conclusion
In this paper, based on Q-learning, we propose a method named straQ*that optimizes long-term returns in emotional support conversation scenarios. Our implementation behaves as a plug-and-play strategic planner which steers the subsequent response generation. We propose two reward mechanisms, straQ*-imit and straQ*-distill, in which the former has higher automatic evaluation results, and the latter performs better in generalization and human preference alignment.
7 Limitation
There are still some limitations of straQ*. The results of the human evaluation may be biased, or deviate from the judgments of actual help-seekers due to the awareness of being engaged in scoring. Then, the test set may be small. Although it has little impact on the comparison between automated and human evaluations, sample sizes for some sub-categories may be insufficient when conducting a detailed analysis.
Appendix A Dataset details
Emotional Support Conversations.
Emotional Support Conversation (ESC) is a task aimed at alleviating users’ negative emotions (e.g., anxiety, depression), where supporters assist seekers in managing emotions triggered by issues like work crises or interpersonal conflicts. Unlike emotion recognition tasks, ESC integrates psychological counseling mechanisms into the dialogue generation process, offering a deeper, context-sensitive solution for emotion regulation. The ESC dataset generally has the following attributes:
  • Emotion: Including emotion types and intensities, which help accurately capture the psychological state of the help-seeker.
  • Help-seeker profile: A brief survey before each conversation that provides insight into the current situation of the help-seeker, revealing the challenges they are facing.
  • Situation: A brief survey before each conversation that provides insight into the current situation of the help-seeker, revealing the challenges they are facing.
  • Strategy: The response rule selected for the current turn based on the seeker’s emotional state. There are eight predefined rules in total.
  • Response: The supporter’s response generated based on the history, inferred state, and selected rule.
In this paper, we mainly study two typical ESC datasets, ESConv and EmpatheticDialogues. ESConv has exactly the aforementioned architecture while EmpatheticDialogues lacks the Strategy. Table 9 summarizes the basic statistical information of ESConv and EmpatheticDialogues.
Category ESConv EmpatheticDialogues
# Sessions 1.3K 2.5K
# Uttr 38K 11.0K
Average # Uttr 28.9 4.3
Average Uttr Len 18.8 16.7
Seeker # Uttr 20K 5.7K
Avg # Uttr 15.4 2.2
Avg Uttr Len 16.8 20.8
# Emotions 11 32
Supporter # Uttr 18K 5.2K
Avg # Uttr 13.6 2.1
Avg Uttr Len 21.0 12.3
# Strategies 8 -
Table 9: Statistics of ESConv and EmpatheticDialogues. ‘Uttr’ abbreviates Utterance.
Emotion Type # Occurrence
anger 111
anxiety 354
depression 334
disgust 40
fear 95
nervousness 13
sadness 308
shame 42
Table 10: Emotion statistics of ESConv.
ESConv.
Motivated by the Helping Skills Theory Hill (2009), Liu et al. (2021) divides ESC into three sequential stages: Exploration, Comforting, and Action, and proposes a dataset called ESConv. For each sample, the conversation is multi-turn, with the dialogue background and user emotion annotated. Upon each utterance of the supporter, 8 distinct support strategies are annotated. Table 11 exhibits an example of ESConv.
Topic I hate my job but I am scared to quit and seek a new career.
Query {history} seeker: Seriously! What I’m scare of now is how to secure another job.
Emotion Anxiety (intensity: 5)
Strategy Reflection of feelings
Response supporter: I can feel your pain just by chatting with you.
Table 11: An example of ESconv.
Strategies Abbr. Definitions
Question Que. Inquiring about problem-related information to help the seeker clarify their issues, using open-ended questions for best results and closed questions for specific details.
Restatement or Paraphrasing Res.& Par. A simple, more concise rephrasing of the help-seeker’s statements that could help them see their situation more clearly.
Reflection of Feelings Ref. Articulate and describe the help-seeker’s feelings.
Self-disclosure Self-Dis. Divulge similar experiences that you have had or emotions that you share with the help-seeker to express your empathy.
Affirmation and Reassurance Aff.& Rea. Affirm the help seeker’s strengths, motivation, and capabilities and provide reassurance and encouragement.
Providing Suggestions Pro. Provide suggestions about how to change, but be careful to not overstep and tell them what to do.
Information Inf. Provide useful information to the help-seeker, for example with data, facts, opinions, resources, or by answering questions.
Others Others Exchange pleasantries and use other support strategies that do not fall into the above categories.
Table 12: Strategy names, abbreviations and detailed definitions in ESConv.
Table 12 provides definitions of support strategies in ESConv. Table 10 lists the emotion types and their occurrences in the dataset. The emotion types include anger, anxiety, depression, disgust, fear, nervousness, sadness, and shame.
EmpatheticDialogues.
EmpatheticDialogues Rashkin et al. (2019) is a dataset that consists of empathetic conversations. It aims to help in the development of empathetic language models by providing a large number of dialogues that express empathy.
Appendix B Detailed Prompts
Instruction template.
To further strengthen the understanding capability of LLM on the strategy selection, we define the instruction as a multi-choice question (MCQ), forcing the LLM to choose one of the option numbers, instead of a plain question. Below is the content of the instruction template $\mathcal{I}$:
You are a psychological consultant providing support to a seeker. The seeker’s basic situation is as follows:
Emotion: {$e$}
Description: {$desp$}
Below is the conversation history between the seeker and the supporter:
{$h$}
The seeker’s current query is:
{$query$}
Based on the above context, please select the most appropriate response strategy from the following options:
strategy #(1) {$a_{1}$}
strategy #(k) {$a_{k}$}
Please provide your selection in the format of (1) through (k). Your selection is:
Generation prompt.
Below is the prompt used by the conversational foundation LLM for the response generation:
You are a psychological consultant providing support to a seeker. The seeker’s basic situation is as follows:
Emotion: {$e$}
Description: {$desp$}
Below is the conversation history between the seeker and the supporter:
{$h$}
The seeker’s current query is:
{$query$}
The current response strategy is:
{$a$}
Based on the current response strategy and other information, please act as a supporter and provide the best response. Keep replies brief without additional pronouns or extra elements.
Prompt of GPT-4 for reward generation.
Below is our prompt of GPT-4 to generate the rewards for straQ*-distill:
You are a psychological consultant providing support to a seeker. The seeker’s basic situation is as follows:
Emotion: {$e$}
Description: {$desp$}
Below is the conversation history between the seeker and the supporter:
{$h$}
The seeker’s current query is:
{$query$}
Please evaluate whether the response is appropriate:
{$resp$}
Based on the information above, evaluate whether the response is suitable. Please remember to respond with a single integer number from 1 to 5, where 1 indicates "not suitable" and 5 indicates "very suitable". Please also provide a brief explanation of your decision.
Table 13: Template of GPT-4 scoring.
Appendix C Principle of Human Scoring
We start with the criteria proposed by Kang et al. (2024). The human evaluation is aimed to align with the ultimate purpose of ESC, the seeker’s satisfaction. To achieve this, the supporter’s behavior can be further classified into the following criteria:
Acceptance: Does the seeker accept without discomfort;
Effectiveness: Is it helpful in shifting negative emotions or attitudes towards a positive direction;
Sensitivity: Does it take into consideration the general state of the seeker. Furthermore, to clarify the capability of LLMs to align strategy and responses, we include Alignment.
To achieve a more elaborate assessment, we consider three more dimensions addressing the generation quality:
Fluency: the level of fluency of response.
Emotion: the emotional intensity of response which could affect the seeker’s emotional state.
Interesting: Whether the response can arouse the seeker’s interest and curiosity, presenting unique ideas, vivid expressions or engaging elements that capture the seeker’s attention and make the interaction more appealing.
We engage our interns as human evaluators to rate the models according to these multiple aspects, namely Fluency, Emotion, Interesting, and Satisfaction, with Satisfaction covering Acceptance, Effective, Sensitivity, and Satisfaction itself.
Throughout this evaluation process, we strictly comply with international regulations and ethical norms, ensuring that all practices conform to the necessary guidelines regarding participant involvement and data integrity.
To guarantee the accuracy and reliability of the evaluation results, a pre-evaluation training program is meticulously designed and implemented. During this training, the evaluation criteria are clearly and systematically expounded. Moreover, detailed explanations and scoring rules corresponding to each score are provided.
Evaluators are required to independently evaluate each sample in strict accordance with the pre-established criteria. By adhering to these principles, the evaluation process maintains objectivity, standardization, and consistency, thus enhancing the overall quality and credibility of the evaluation results.
The detailed manual scoring criteria are as follows:
  • Fluency:
    1: The sentence is highly incoherent, making it extremely difficult to understand and failing to convey a meaningful idea.
    2: The sentence has significant incoherence issues, with only parts of it making sense and struggling to form a complete thought.
    3: The sentence contains some incoherence and occasional errors, but can still convey the general meaning to a certain extent.
    4: The sentence is mostly fluent with only minor errors or slight awkwardness in expression, and effectively communicates the intended meaning.
    5: Perfect. The sentence is completely fluent, free of any errors in grammar, punctuation, or expression, and clearly conveys the idea.
  • Emotion:
    1: The emotional expression is extremely inappropriate and chaotic, not in line with the content, and may convey wrong emotions.
    2: The emotional expression has obvious flaws, either too weak or exaggerated, and is disjointed from the content.
    3: The emotional expression is average. It can convey basic emotions but lacks depth and has minor issues.
    4: The emotional expression is good. It can effectively convey the intended emotion with an appropriate intensity and is well integrated with the content.
    5: The emotional expression is excellent. It is rich, nuanced, and perfectly matches the content, capable of evoking a strong and appropriate emotional response.
  • Acceptance:
    1: The response inescapably triggers emotional resistance.
    2: The response is highly likely to trigger emotional resistance.
    3: The response has a possibility of emotional resistance occurring.
    4: The response rarely provokes emotional resistance.
    5: The response has no occurrence of emotional resistance.
  • Effectiveness:
    1: The response actually worsens the seeker’s emotional distress.
    2: The response carries the risk of increasing stress levels, and this outcome varies depending on the individual user.
    3: The response fails to alter the seeker’s current emotional intensity and keeps it at the same level.
    4: The response shows promise in calming the emotional intensity; however, it is overly complicated or ambiguous for the user to fully comprehend and utilize effectively.
    5: The response appears to be highly effective in soothing the seeker’s emotions and offers valuable and practical emotional support.
  • Sensitivity:
    1: The response renders inaccurate evaluations regarding the seeker’s state.
    2: The response is characterized by rash judgments, as it lacks adequate assessment and in-depth exploration of the seeker’s state.
    3: The response is formulated with a one-sided judgment and a limited exploration of the seeker’s state.
    4: The response demonstrates an understanding that only covers a part of the seeker’s state.
    5: The response precisely grasps the seeker’s state and is appropriately tailored according to the seeker’s actual situation.
    Strategy Count Ratio Max Min Avg Median
    Que. 2574 17.6% 5 1 3.54 4
    Res.& Par. 981 6.7% 5 1 3.48 3
    Ref. 1253 8.6% 5 2 3.65 4
    Self-Dis. 1410 9.6% 5 2 3.48 3
    Aff.& Rea. 2444 16.7% 5 1 3.76 4
    Pro. 2367 16.2% 5 1 3.77 4
    Inf. 995 6.8% 5 2 3.75 4
    Others 2600 17.8% 5 1 3.77 4
    Total 14624 100.0% 5 1 3.67 4
    Table 14: Statistics of GPT-4 score.
  • Alignment:
    1: The response is in total contradiction to the predicted strategy.
    2: The response has a minor deviation from the predicted strategy.
    3: There is some ambiguity between the response and the predicted strategy.
    4: The response largely matches the predicted strategy, yet it contains some ambiguous elements.
    5: The response effectively makes itself consistent with the predicted strategy.
  • Satisfaction:
    1: The response is extremely disappointing. It doesn’t answer the question at all and is of no help.
    2: The response is poor. It only gives a partial answer and leaves many doubts unresolved.
    3: The response is average. It meets the basic requirements but isn’t particularly outstanding.
    4: The response is good. It answers the question clearly and provides some useful details.
    5: The response is excellent. It not only answers the question perfectly but also offers valuable additional insights.
Appendix D Details of GPT-4 Scoring
Table 14 presents GPT-4 score statistics across different response strategies. The overall average score is 3.67, with a median of 4. The most frequently used strategies are Others (17.8%), Questioning (17.6%), and Affirmation & Reasoning (16.7%), while Restating & Paraphrasing (6.7%) and Information Providing (6.8%) appear less often. In terms of average score, Providing Opinions, Others, and Affirmation & Reasoning score the highest (all-around 3.76–3.77), whereas Restating & Paraphrasing and Self-Disclosure have the lowest average scores (3.48).